lil (6) Let L in R3 be the line through the origin spanned by the vector v= 1. Find the linear equations that define L, i.e.,Let L in R 3 be the line through the origin spanned by the vector v =   1 1 3  . Find the linear equations that define L, i.e., find a system of linear equations whose solutions are the points in L.
Answer #1

The equation of L is,

(x-0)/1 = (y-0)/1 = (z-0)/3

i.e. x/1 = y/1 = z/3

Or, x = y = z/3